OpenStudy (anonymous):

Write the algebraic equation of: Five-sixths of a number is -60. What is the number?

OpenStudy (anonymous):

\[(5\div60)x =-60\]

OpenStudy (anonymous):

Five sixths of a number means: Let the number be x.. So Five sixths of this number x will be : \(\frac{5}{6} \; of \;x = \frac{5}{6}x\)

OpenStudy (anonymous):

This is given equal to -60.. So: \[\frac{5}{6}x = -60\] Now here you have to find x..

OpenStudy (anonymous):

Getting this much @rannsan

OpenStudy (anonymous):

I am trying to follow

OpenStudy (anonymous):

Read the question and tell me what is that you are not understanding?? Is question clear to you or not ??

OpenStudy (anonymous):

it is clear I am following you I am just confused as to how to find what x is.

OpenStudy (anonymous):

You have to get rid of 5/6 there.. Can you multiply both the sides by 6??

OpenStudy (anonymous):

just multiply -60 with 6 and divide by 5

OpenStudy (anonymous):

\[\frac{5}{6}x \times 6= -60 \times 6\]

OpenStudy (anonymous):

(-60)6 and divide by 5?

OpenStudy (anonymous):

Can you solve that I have written above ??

OpenStudy (anonymous):

NO

OpenStudy (anonymous):

Can't you cancel 6 with 6 on Left hand side ??

OpenStudy (anonymous):

5X = -60?

OpenStudy (anonymous):

No.. You have to multiply 6 with -60.. On Right hand side why you cancelled 6 ??

OpenStudy (anonymous):

oops I misunderstood that one

OpenStudy (anonymous):

\[\frac{5x}{\cancel{6}} \times \cancel{6} = -60 \times 6\] \[5x = -360\] Now divide by 5 both the sides to get x..

OpenStudy (anonymous):

\[5x \times = -60 \times 5x = -360

OpenStudy (anonymous):

I am still not following

OpenStudy (anonymous):

Where you lost ??

OpenStudy (anonymous):

on everything...

OpenStudy (anonymous):

I need step by step once again please.

OpenStudy (anonymous):

\[\frac{5}{6}x \times 6= -60 \times 6\] Till this are you clear ??

OpenStudy (anonymous):

(5/6x)6 = -60

OpenStudy (anonymous):

Why are not multiplying 6 on right hand side ??

OpenStudy (anonymous):

I thought I was

OpenStudy (anonymous):

Multiplying 6 both sides means you have to multiply 6 on left as well as right hand side.. You have multiplied with left hand side only..

OpenStudy (anonymous):

oh ok let me see

OpenStudy (anonymous):

(5/6x)6 = -60x6

OpenStudy (anonymous):

Yes..

OpenStudy (anonymous):

Now on left hand side you can cancel 6 with bottom 6 right ??

OpenStudy (anonymous):

(5x)6 = -60x6

OpenStudy (anonymous):

Do you know how we cancel things ??

OpenStudy (anonymous):

no

OpenStudy (anonymous):

See: you have cancel one 6 in left hand side.. but how can you cancel one 6?? you can cancel two 6 and that too will be in the form: 6 - 6 you can cancel them. 6/6 here also you can cancel them.. Getting ??

OpenStudy (anonymous):

See like this: \[\frac{5x}{\cancel{6}} \times \cancel{6} = -60 \times 6\] So you are left with now: \[5x = -360\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Sure no?? Getting ??

OpenStudy (anonymous):

I am getting it, writing notes

OpenStudy (anonymous):

Now you are just one step away: You just need to divide by 5 both the sides to get rid of 5 attached to x.. can you do that? Both the sides means you will be divide left as well as right hand side by 5.. Show me what you did..

OpenStudy (anonymous):

Are you really typing @rannsan

OpenStudy (anonymous):

5(5x) = (5)-360 25= -360(5) 25=-503.8 Ok I know that is wrong

OpenStudy (anonymous):

I think you are multiplying.. I said to divide.. You know what divide means ??

OpenStudy (anonymous):

You did say divide I was multiplying

OpenStudy (anonymous):

\[5\div5x = 1\]

OpenStudy (anonymous):

\[-360\div5\]

OpenStudy (anonymous):

No you are still doing opposite: \[\frac{5x}{5} = \frac{-360}{5}\] Can you do this now ??

OpenStudy (anonymous):

5x/5 = -360/5

OpenStudy (anonymous):

Yes..

OpenStudy (anonymous):

Now when you cancel top 5 with bottom 5 you will left with x and not 1.. Right ??

OpenStudy (anonymous):

x only

OpenStudy (anonymous):

Yes.. This is on left hand side.. now can you solve right hand side??

OpenStudy (anonymous):

x= -360/5

OpenStudy (anonymous):

I divide the 5 by the -360?

OpenStudy (anonymous):

Yes.. go ahead..

OpenStudy (anonymous):

\[\frac{-360}{5} = ??\]

OpenStudy (anonymous):

-647.8

OpenStudy (anonymous):

I know ight isn't right

OpenStudy (anonymous):

-360/5

OpenStudy (anonymous):

\[\frac{-360}{5} = -72\]

OpenStudy (anonymous):

-360/5=-72

OpenStudy (anonymous):

thank you! once again